Suppose that three random variables X1,X2,X3 form a random sample from a
distribution for which the mean is 5. Determine the value of E(2X1 - 3X2 +X3 - 4).
1
Expert's answer
2019-11-05T09:32:39-0500

E(2X1-3X2+X3-4)=E(2X1)- E(3X2)+E(X3) -4 =2E(X1)- 3E(X2)+E(X3) -4=2*5-3*5+5-4=10 - 15+ 5 -4= -4